# Model Merging in LLMs, MLLMs, and Beyond: Methods, Theories, Applications and Opportunities

### v3

@@ 4.3.2 Model Merging for Better Domain Generalization
Domain generalization methods aim to generalize to an unknown target domain using only training data from source domains. For instance, in the context of traffic sign recognition, the training data for a machine learning (ML) model tasked with identifying traffic signs in various urban environments come from multiple cities (i.e., source domains). However, when deployed, the model must recognize traffic signs in new urban environments (i.e., target domains) that it has never encountered before. Existing DG methods can be classified into domain alignment, data augmentation, regularization, and meta-learning frameworks [8]. Complementary to these approaches, model merging techniques can be seamlessly integrated to further improve out-of-domain performance without modification. Specifically, model merging in DG mainly occurs during the training process of the source domain model. Merging the intermediate weight states from different training stages helps improve the stability and generalization of the final model.
@@ 4.5 Model Merging in Zero-shot/Few-shot Learning
In practical applications of machine learning models, collecting a large amount of labeled data can be expensive or infeasible in specific scenarios (e.g., medical diagnosis, real-time monitoring). Users often want deep models to effectively perform new tasks that have not been encountered before, that is, an ability commonly referred to as cross-task generalization [61]. Zero-shot [115] and few-shot learning [198] can reduce the dependence on large amounts of data and allow the model to better deal with unseen categories or small numbers of samples, improving the cross-task generalization ability of the model. In few-shot learning, a common approach is to fine-tune the model using the limited examples available. However, because of the minimal data, this fine-tuning process is often unstable and yields only modest performance improvements. Recently, some studies have explored merging pre-trained models (from some publicly accessible resources) to enhance cross-task generalization under zero-shot and few-shot conditions.
@@ 4.5.2 Model Merging for Cross-task Generalization in Few-shot Learning
Recent work on model merging suggests a positive answer, showing that merged models can enhance generalization in few-shot settings [53, 9, 61]. For example, LoraHub [61] proposes to merge LoRA modules available on HuggingFace to achieve adaptive performance for unseen tasks, where the merging coefficients of different LoRA are searched in a black-box gradient-free manner with few-shot samples. As expected, few-shot LoraHub performs better than few-shot in-context learning and reduces inference costs by eliminating the need for examples as input to LLMs. LoraRetriever [212] further proposes dynamically retrieving the most relevant LoRAs based on the input and merging them. Similarly, MerA [53] proposes merging pretrained adapters into a single adapter for few-shot NLP scenarios. In general, well-trained LoRAs or adapters can serve as valuable resources that users can easily share, access, and apply to a variety of downstream tasks. In the real world, upstream and downstream tasks can be entirely disparate, originating from different datasets, domains, or even different parts of the same dataset. Asadi et al. [9] comprehensively evaluates model merging in the few-shot learning setting. Specifically, this study examines three cases of label, domain, and task drift between upstream and downstream tasks. The results demonstrate that model merging enhances the model’s generalization ability in few-shot learning scenarios across different contexts.

### v4

@@ 4.3.2 Model Merging for Better Domain Generalization
Domain generalization methods aim to generalize to an unknown target domain using only training data from source domains. For instance, in the context of traffic sign recognition, the training data for a machine learning (ML) model tasked with identifying traffic signs in various urban environments come from multiple cities (i.e., source domains). However, when deployed, the model must recognize traffic signs in new urban environments (i.e., target domains) that it has never encountered before. Existing DG methods can be classified into domain alignment, data augmentation, regularization, and meta-learning frameworks [8]. Complementary to these approaches, model merging techniques can be seamlessly integrated to further improve out-of-domain performance without modification [17, 8, 109]. Specifically, model merging in DG mainly occurs during the training process of the source domain model. Merging the intermediate weight states from different training stages helps improve the stability and generalization of the final model.
@@ 4.5 Model Merging in Zero-shot/Few-shot Learning
In practical applications of machine learning models, collecting a large amount of labeled data can be expensive or infeasible in specific scenarios (e.g., medical diagnosis, real-time monitoring). Users often want deep models to effectively perform new tasks that have not been encountered before, that is, an ability commonly referred to as cross-task generalization [67]. Zero-shot [128] and few-shot learning [224] can reduce the dependence on large amounts of data and allow the model to better deal with unseen categories or small numbers of samples, improving the cross-task generalization ability of the model. In few-shot learning, a common approach is to fine-tune the model using the limited examples available. However, because of the minimal data, this fine-tuning process is often unstable and yields only modest performance improvements. Recently, some studies have explored merging pre-trained models (from some publicly accessible resources) to enhance cross-task generalization under zero-shot and few-shot conditions [211, 67, 79, 137, 129, 196, 181].
@@ 4.5.2 Model Merging for Cross-task Generalization in Few-shot Learning
Recent work on model merging suggests a positive answer, showing that merged models can enhance generalization in few-shot settings [58, 9, 67, 196]. For example, LoraHub [67] proposes to merge LoRA modules available on HuggingFace to achieve adaptive performance for unseen tasks, where the merging coefficients of different LoRA are searched in a black-box gradient-free manner with few-shot samples. As expected, few-shot LoraHub performs better than few-shot in-context learning and reduces inference costs by eliminating the need for examples as input to LLMs. LoraRetriever [237] further proposes dynamically retrieving the most relevant LoRAs based on the input and merging them. Similarly, MerA [58] proposes merging pretrained adapters into a single adapter for few-shot NLP scenarios. In general, well-trained LoRAs or adapters can serve as valuable resources that users can easily share, access, and apply to a variety of downstream tasks. In the real world, upstream and downstream tasks can be entirely disparate, originating from different datasets, domains, or even different parts of the same dataset. Asadi et al. [9] comprehensively evaluates model merging in the few-shot learning setting. Specifically, this study examines three cases of label, domain, and task drift between upstream and downstream tasks. The results demonstrate that model merging enhances the model’s generalization ability in few-shot learning scenarios across different contexts.